# An Introduction to Single-Antenna Radio Astronomical Polarimetry

### v2

@@ 6.1 Phase convention
6.2 Receiver gains
The receiver gains matrix, G, is often determined using observations of an artificial reference source with a well-defined polarization state. For example, many receivers have a built-in source of broadband noise (e.g., a noise diode coupled to the receptors) that can be used to estimate gains as a function of radio frequency. Determination of G using only an artificial noise source is based on the ideal feed assumptions, in which the receptors are orthogonal and there is no cross-coupling between them (i.e., D is the identity matrix); and the artificial reference source illuminates the receptors equally (i.e., the receptors have identical responses to the reference source, in both amplitude and phase). Although these assumptions are common, it is more accurate to experimentally determine both D and the Stokes parameters intrinsic to the artificial reference source (e.g., van Straten, 2004), as described in Section 6.3.
Even when fully modeling the deviations from an ideal feed and polarization of the artificial reference source, the ideal feed assumptions can provide a useful, and sometimes necessary, first guess for G. Therefore, they are an important part of polarimetric calibration for both single antennas (e.g., von Hoensbroech & Xilouris, 1997; Navarro et al., 1997; Gould & Lyne, 1998; Han et al., 2009) and interferometric arrays (e.g., Johnston et al., 2008; Serylak et al., 2021). As shown in Section 5.2, the polar decomposition of the complex gains,
@@ 6.2 Receiver gains
$\underline{\underline{\mbox{$S$}}}_{\mathrm{ref}}=I_{\mathrm{ref}}[1,0,1,0],$ (92)
@@ 6.2.1 Differential Phase
6.2.2 Differential Gain
As a boost along the $S_{1}$ axis, differential gain mixes only $S_{0}$ and $S_{1}$; therefore, determination of $\gamma$ requires observations of a polarized reference source with known intrinsic values of $S_{0}$ and $S_{1}$, where $S_{1}$ corresponds to Stokes Q in the linear basis and to Stokes V in the circular basis. Given the measured Stokes parameters of the reference source, with observed values of $S_{0}^{\prime}$ and $S_{1}^{\prime}$,, equations (17) and (18) can be solved for the intrinsic and observed values of $\langle|e_{0}(t)|^{2}\rangle$ and $\langle|e_{1}(t)|^{2}\rangle$; e.g.
@@ 6.2.2 Differential Gain
6.3 Deviations from ideal
For a single antenna, the deviations from an ideal feed, D, can be determined by modeling variations in observed polarization of an astrophysical source as a function of parallactic angle (e.g., Stinebring et al., 1984; Xilouris, 1991; McKinnon, 1992; Heiles et al., 2001; Johnston, 2002; van Straten, 2004). In some treatments, D is estimated after the complex receiver gains, G, have been calibrated; in others, D and G are jointly determined. Some of the models used to describe $\bf D$} are discussed and compared in Section 7. In an interferometer, unique values of ${\bf J}_{i}={\bf G}_{i}{\bf D}_{i}$ must be determined for each element in the array, indexed by $i$. When used as a phased-array, only the Jones matrix of the sum, ${\bf J}=\sum_{i}{\bf J}_{i},$ must be determined.
@@ 6.3 Deviations from ideal
In either case (free-space or guided transmission), the instrumental response to an artificial noise source typically differs from its response to an astrophysical source. However, in many experiments, the absolute polarization of the artificial reference source is irrelevant. For example, techniques used to update the differential gain and phase given known values of $\mbox{$\rho$}_{\mathrm{ref}}$ and D (e.g., Ord et al., 2004) rely only on the temporal stability of these quantities. Therefore any differences in the instrumental response to astrophysical and artificial sources can be absorbed by redefinition of $\mbox{$\rho$}_{\mathrm{ref}}$.
@@
6.4 Nominal Feed Configuration
@@
6.5 Celestial sphere projection
@@
where $\theta_{\mathrm{feed}}$ is the feed horn rotation, ${\bf J}(l,m)$ models direction-dependent effects of the antenna (e.g., as a function of the direction cosines $l$ and $m$ that describe angular offsets from the primary axis), and $\theta_{\mathrm{para}}$ is the parallactic angle.
6.6 Faraday Rotation
@@
7 Polarimeter Model Selection
@@ 7.3 Injectivity
Section 5.1 of H2001 describes ambiguity in the ellipticity angle, $\alpha$, and differential phase, $\psi$, that arises in the case of linearly-polarized receptors. They assert that two possible solutions, $(\alpha_{1},\psi_{1})=(0,\psi_{0})$ and $(\alpha_{2},\psi_{2})=(\pi/2,\psi_{0}+\pi)$, differ by only a 180 rotation about the Stokes V axis, which is assumed to be zero and therefore results in negation of the unknown Stokes Q and U of the calibrator source. Referring to Appendix D.3 and Section 4 of B2000, this ambiguity can be eliminated by restricting the ellipticity angle to the interval $\pi/4\leq\alpha\leq\pi/4$.
7.4 Self-consistency
@@
Consider the observation of an ideal reference source (eqn. 92) with Stokes parameters, $\underline{\underline{\mbox{$S$}}}=I[1,0,1,0]$. The transformed Stokes parameters,
$\underline{\underline{\mbox{$S$}}}^{\prime}={\underline{\underline{{\bf M}}}}_{A}\underline{\underline{\mbox{$S$}}}=I[1,\Delta G/2,\cos\psi,\sin\psi],$ (108)
@@
As no linear transformation of the electric field vector can convert a valid polarization state into an over-polarized state (Proof 4.3), ${\underline{\underline{{\bf M}}}}_{A}$ cannot have an equivalent Jones matrix and therefore must be impure. Not all impure Muller matrices result in over-polarization; some impure transformations depolarize a purely polarized state (e.g. see Appendix C.7), and no linear transformation of the electric field is able to do so (Proof 4.4). This specific example demonstrates the potential pitfalls of using Mueller matrices and small-value approximations during numerical analysis.
7.5 Numerical Stability
@@ 7.5.1 Fundamental Instability
The analysis of fundamental instability is extended in Appendix I, which identifies additional degenerate dof in impure Mueller matrices, owing to unknown components of the instrumental response that commute with the constraining transformation. It also presents an example of parameter collinearity that arises when an unknown component of the Stokes parameters intrinsic to the observed sources is an eigenvector of the constraining transformation. Owing to the degenerate sign of Stokes V identified in Appendix I.1, the phase convention of the backend cannot be determined using only unknown sources observed at multiple parallactic angles, regardless of the basis defined by the polarization of the receptors.
@@ 7.5.3 Intrinsic Instability
8 Conclusion
The geometry of the polarization ellipse that defines the Stokes parameters is elegantly connected to that of the coherency matrix via the Pauli matrices. Geometry also relates linear transformations of the electric field to their impact on the Stokes parameters, leading to a physically meaningful and broadly applicable classification of polarimetric transformations. The properties of these transformations simplify the analysis of the elements of the signal path and yield practical guidelines that can be applied during calibration.
@@ 8 Conclusion
Current and future scientific goals, such as detecting the neutral Hydrogen emission from the epoch of reionization (e.g., Trott et al., 2018; Mertens et al., 2020) and understanding the origin and evolution of interstellar and intergalactic magnetic fields (e.g., Han, 2017), motivate the development of next-generation facilities and methods of polarimetric analysis, such as Faraday tomography (e.g., Ganguly et al., 2000; Van Eck et al., 2017). Furthermore, the unprecedented sensitivity, resolution, and fields of view of new observatories necessitate further research and development of novel methods of direction-dependent interferometric calibration (e.g., Smirnov & Tasse, 2015) and correction of ionospheric Faraday rotation (e.g., de Gasperin et al., 2019). Improved polarimetric fidelity also has the potential to increase the sensitivity of pulsar timing array experiments to the low-frequency stochastic gravitational wave background (e.g., Guillemot et al., 2023; Rogers et al., 2024).
@@
I’m grateful to the mentors who sparked my interest in polarimetry, and the colleagues who share their enthusiasm for the subject, especially Matthew Britton, Paul Demorest, Russell Edwards, Marisa Geyer, Johan Hamaker, JinLin Han, Simon Johnston, Aris Karastergiou, Vivek Venkatraman Krishnan, Dick Manchester, Maaijke Mevius, Aris Noutsos, Steve Ord, Stefan Osłowski, Aditya Parthasarathy, John Reynolds, Maciej Serylak, Ben Stappers, and Caterina Tiburzi. Marisa’s visit in 2018 prompted the development of this tutorial, and Caterina and Maaijke provided insightful feedback and brilliant ideas that significantly enhanced its elaboration. The comprehensive and constructive criticism of the anonymous reviewer led to further additions and improvements.
@@ Appendix B Vector, Matrix, and Tensor Notation
$\mbox{$r$}=\left(r^{x}\;\;r^{y}\right).$ (B2)
@@
$\mbox{$e$}^{\dagger}=\left(e_{x}^{*}\;\;e_{y}^{*}\right).$ (B3)
@@
A double contraction of two tensors, $A:B$, is defined as a tensor product followed by summation over two pairs of matching indeces, thereby reducing the rank of the result of the tensor product by four. For example, the double contraction of a rank 4 tensor $\mathcal{U}$ with a (rank 2) matrix C yields another (rank 2) matrix with elements given by
$\left\{\mathcal{U}:{\bf C}\right\}_{\mu}^{\nu}=\mathcal{U}_{\mu\kappa}^{\nu\lambda}C_{\lambda}^{\kappa}.$ (B13)
@@
${\rm Tr}\left[{{\bf A}{\bf B}}\right]={\rm Tr}\left[{{\bf B}{\bf A}}\right]=A_{\beta}^{\gamma}B_{\gamma}^{\beta}={\bf A}:{\bf B}={\bf B}:{\bf A}.$ (B15)
That is, the trace of a matrix product is equivalent to the double contraction of the matrices, also known as the projection with respect to the trace inner product. Like the scalar product of two vectors, this projection is symmetric and matrices A and B are orthogonal with respect to the trace inner product if ${\bf A}:{\bf B}=0$.
@@
$\left\{\left({\bf A}\otimes{\bf B}\right):{\bf C}\right\}_{\mu}^{\nu}$ $=\left\{{\bf A}\otimes{\bf B}\right\}_{\mu\kappa}^{\nu\lambda}C_{\lambda}^{\kappa}$ Eqn. (B13) $=A_{\mu}^{\nu}B_{\kappa}^{\lambda}C_{\lambda}^{\kappa}$ Eqn. (B10) $=A_{\mu}^{\nu}({\bf B}:{\bf C})$ Eqn. (B15) $=\left\{{\bf A}\left({\bf B}\,\mbox{$:$}\,{\bf C}\right)\right\}_{\mu}^{\nu}$
@@
$\left\{\left({\bf A}\,\tilde{\otimes}\,{\bf B}\right):{\bf C}\right\}_{\mu}^{\nu}$ $=\left\{{\bf A}\,\tilde{\otimes}\,{\bf B}\right\}_{\mu\kappa}^{\nu\lambda}C_{\lambda}^{\kappa}$ Eqn. (B13) $=A_{\mu}^{\lambda}B_{\kappa}^{\nu}C_{\lambda}^{\kappa}$ Eqn. (45) $=A_{\mu}^{\lambda}\left\{{\bf C}\cdot{\bf B}\right\}_{\lambda}^{\nu}$ Eqn. (B7) $=\left\{{\bf A}{\bf C}{\bf B}\right\}_{\mu}^{\nu}$ Eqn. (B7)
@@ C.1 Hermitian Basis Matrices
$\mbox{$\sigma$}_{\mu}:\mbox{$\sigma$}_{\nu}$ $=$ $2\delta_{\mu\nu}$ (C4)
@@ C.2 Basis Transformations
In general, a unitary matrix has a determinant equal to $\exp(i\phi)$, and multiplication by $\exp(-i\phi/2)$ yields a unimodular matrix (with $|{\bf R}|=1$). Basis transformations defined by unitary matrices are used extensively throughout both Section 5 and the following sections of the appendix.
@@ C.3 Spectral Decomposition
Both Hermitian and unitary matrices are normal, and each takes diagonal form in the natural basis defined by its eigenvectors. Consequently, any Hermitian matrix can be diagonalized such that it is equivalent to a differential gain transformation (see Section 5.2.1) in its natural basis. Similarly, any unitary matrix is equivalent to differential phase (see Section 5.2.2) in its natural basis. This is applied in Proof 5.2, where rotation about the line of sight is shown to be equivalent to differential phase between the circularly-polarized natural modes of a magnetized cold plasma.

### v4

@@ VI.1 Phase Convention
VI.2 Receiver Gains
The receiver gains matrix, G, is often determined using observations of an artificial reference source with a well-defined polarization state. For example, many receivers have a built-in source of broadband noise (e.g., a noise diode coupled to the receptors) that can be used to estimate gains as a function of radio frequency. Determination of G using only an artificial noise source is based on the ideal feed assumptions, in which the receptors are orthogonal and there is no cross-coupling between them (i.e., D is the identity matrix); and the artificial reference source illuminates the receptors equally (i.e., the receptors have identical responses to the reference source, in both amplitude and phase). Although these assumptions are common, it is more accurate to experimentally determine both D and the Stokes parameters intrinsic to the artificial reference source (e.g., van Straten, 2004), as described in Section VI.3.
Even when fully modeling the deviations from an ideal feed and polarization of the artificial reference source, the ideal feed assumptions can provide a useful, and sometimes necessary, first guess for G. Therefore, they are an important part of polarimetric calibration for both single antennas (e.g., von Hoensbroech & Xilouris, 1997; Navarro et al., 1997; Gould & Lyne, 1998; Han et al., 2009) and interferometric arrays (e.g., Johnston et al., 2008; Serylak et al., 2021). As shown in Section V.2, the polar decomposition of the complex gains,
@@ VI.2 Receiver Gains
$\underline{\underline{\mbox{$S$}}}_{\mathrm{ref}}=I_{\mathrm{ref}}[1,0,1,0]^{T},$ (92)
@@ VI.2.1 Differential Phase
VI.2.2 Differential Gain
As a boost along the $S_{1}$ axis, differential gain mixes only $S_{0}$ and $S_{1}$; therefore, determination of $\gamma$ requires observations of a polarized reference source with known intrinsic values of $S_{0}$ and $S_{1}$, where $S_{1}$ corresponds to Stokes Q in the linear basis and to Stokes V in the circular basis. Given the measured Stokes parameters of the reference source, with observed values of $S_{0}^{\prime}$ and $S_{1}^{\prime}$, Equations (17) and (18) can be solved for the intrinsic and observed values of $\langle|e_{0}(t)|^{2}\rangle$ and $\langle|e_{1}(t)|^{2}\rangle$; e.g.
@@ VI.2.2 Differential Gain
VI.3 Deviations from Ideal
For a single antenna, the deviations from an ideal feed, D, can be determined by modeling variations in the observed polarization of an astrophysical source as a function of parallactic angle (e.g., Stinebring et al., 1984; Xilouris, 1991; McKinnon, 1992; Heiles et al., 2001; Johnston, 2002; van Straten, 2004). In some treatments, D is estimated after the complex receiver gains, G, have been calibrated; in others, D and G are jointly determined. Some of the models used to describe $\bf D$} are discussed and compared in Section VII. In an interferometer, unique values of ${\bf J}_{i}={\bf G}_{i}{\bf D}_{i}$ must be determined for each element in the array, indexed by $i$. When used as a phased-array, only the Jones matrix of the sum, ${\bf J}=\sum_{i}{\bf J}_{i},$ must be determined.
@@ VI.3 Deviations from Ideal
In either case (free-space or guided transmission), the instrumental response to an artificial noise source typically differs from its response to an astrophysical source. However, in many experiments, the absolute polarization of the artificial reference source is irrelevant. For example, techniques used to update the differential gain and phase given known values of $\mbox{$\rho$}_{\mathrm{ref}}$ and D (e.g., Ord et al., 2004) rely only on the temporal stability of these quantities. Therefore, any differences in the instrumental response to astrophysical and artificial sources can be absorbed by redefinition of $\mbox{$\rho$}_{\mathrm{ref}}$.
@@
VI.4 Nominal Feed Configuration
@@
VI.5 Celestial Sphere Projection
@@
where $\theta_{\mathrm{feed}}$ is the feed horn rotation, ${\bf J}(l,m)$ models direction-dependent effects of the antenna (e.g., as a function of the direction cosines $l$ and $m$ that describe angular offsets from the primary axis), and $\theta_{\mathrm{para}}$ is the parallactic angle (e.g., Guillemot et al., 2025).
VI.6 Faraday Rotation
@@
VII Polarimeter Model Selection
@@ VII.3 Injectivity
Section 5.1 of H2001 describes ambiguity in the ellipticity angle, $\alpha$, and differential phase, $\psi$, that arises in the case of linearly-polarized receptors. They assert that two possible solutions, $(\alpha_{1},\psi_{1})=(0,\psi_{0})$ and $(\alpha_{2},\psi_{2})=(\pi/2,\psi_{0}+\pi)$, differ by only a 180 rotation about the Stokes V axis, which negates the unknown values of Stokes Q and U of the calibrator source. Referring to Appendix D.3 and Section 4 of B2000, this ambiguity can be eliminated by restricting the ellipticity angle to the interval $\pi/4\leq\alpha\leq\pi/4$.
VII.4 Self-consistency
@@
Consider the observation of an ideal reference source (Eqn. 92) with Stokes parameters, $\underline{\underline{\mbox{$S$}}}=I[1,0,1,0]^{T}$. The transformed Stokes parameters,
$\underline{\underline{\mbox{$S$}}}^{\prime}={\underline{\underline{{\bf M}}}}_{A}\underline{\underline{\mbox{$S$}}}=I[1,\Delta G/2,\cos\psi,\sin\psi]^{T},$ (108)
@@
As no linear transformation of the electric field vector can convert a valid polarization state into an over-polarized state (Proof IV.3), ${\underline{\underline{{\bf M}}}}_{A}$ cannot have an equivalent Jones matrix and therefore must be impure. Not all impure Muller matrices result in over-polarization; some impure transformations depolarize a purely polarized state (e.g. see Appendix C.7), and no linear transformation of the electric field is able to do so (Proof IV.4). This specific example demonstrates the potential pitfalls of using Mueller matrices and small-value approximations during numerical analysis.
VII.5 Numerical Stability
@@ VII.5.1 Fundamental Instability
The analysis of fundamental instability is extended in Appendix I, which identifies additional degenerate dof in impure Mueller matrices, owing to unknown components of the instrumental response that commute with the matrix argument. It also presents an example of parameter collinearity that arises when an unknown component of the Stokes parameters intrinsic to the observed sources is an eigenvector of the matrix argument. Owing to the degenerate sign of Stokes V identified in Appendix I.1, the phase convention of the backend cannot be determined using only unknown sources observed at multiple parallactic angles, regardless of the basis defined by the polarization of the receptors.
@@ VII.5.3 Intrinsic Instability
VIII Conclusion
The geometry of the polarization ellipse that defines the Stokes parameters is elegantly connected to that of the coherency matrix via the Pauli matrices. Geometry also relates linear transformations of the electric field to their impact on the Stokes parameters, leading to a physically meaningful and broadly applicable classification of polarimetric transformations. The properties of these transformations simplify the analysis of the elements of the signal path and yield both theoretical insights and practical guidelines that can be applied during calibration.
@@ VIII Conclusion
Current and future scientific goals, such as detecting the neutral Hydrogen emission from the epoch of reionization (e.g., Trott et al., 2018; Mertens et al., 2020) and understanding the origin and evolution of interstellar and intergalactic magnetic fields (e.g., Han, 2017), motivate the development of next-generation facilities and methods of polarimetric analysis, such as Faraday tomography (e.g., Ganguly et al., 2000; Van Eck et al., 2017). Furthermore, the unprecedented sensitivity, resolution, and fields of view of new observatories necessitate further research and development of novel methods of direction-dependent interferometric calibration (e.g., Smirnov & Tasse, 2015) and correction of ionospheric Faraday rotation (e.g., de Gasperin et al., 2019). Improved polarimetric fidelity also has the potential to increase the sensitivity of pulsar timing array experiments to the low-frequency stochastic gravitational wave background (e.g., Guillemot et al., 2023; Rogers et al., 2024; Guillemot et al., 2025).
@@
I’m grateful to the mentors who sparked my interest in polarimetry and the colleagues who share their enthusiasm for the subject, especially Matthew Britton, Paul Demorest, Russell Edwards, Marisa Geyer, Lucas Guillemot, Johan Hamaker, JinLin Han, Simon Johnston, Aris Karastergiou, Vivek Venkatraman Krishnan, Dick Manchester, Maaijke Mevius, Aris Noutsos, Steve Ord, Stefan Osłowski, Aditya Parthasarathy, John Reynolds, Maciej Serylak, Ben Stappers, and Caterina Tiburzi. Marisa’s visit in 2018 prompted the development of this tutorial, and Caterina and Maaijke provided insightful feedback and brilliant ideas that significantly enhanced its elaboration. The comprehensive and constructive criticism of the anonymous reviewer led to further additions and improvements.
@@ Appendix B Vector, Matrix, and Tensor Notation
$\mbox{$r$}=\left(r^{x},\,r^{y}\right).$ (B2)
@@
$\mbox{$e$}^{\dagger}=\left(e_{x}^{*},\,e_{y}^{*}\right).$ (B3)
@@
A double contraction of two tensors, $A\mathbin{:}B$, is defined as a tensor product followed by summation over two pairs of matching indeces, thereby reducing the rank of the result of the tensor product by four. For example, the double contraction of a rank 4 tensor $\mathcal{U}$ with a (rank 2) matrix C yields another (rank 2) matrix with elements given by
$\left\{\mathcal{U}\mathbin{:}{\bf C}\right\}_{\mu}^{\nu}=\mathcal{U}_{\mu\kappa}^{\nu\lambda}C_{\lambda}^{\kappa}.$ (B13)
@@
${\rm Tr}\left[{{\bf A}{\bf B}}\right]={\rm Tr}\left[{{\bf B}{\bf A}}\right]=A_{\beta}^{\gamma}B_{\gamma}^{\beta}={\bf A}\mathbin{:}{\bf B}={\bf B}\mathbin{:}{\bf A}.$ (B15)
That is, the trace of a matrix product is equivalent to the double contraction of the matrices, also known as the projection with respect to the trace inner product. Like the scalar product of two vectors, this projection is symmetric and matrices A and B are orthogonal with respect to the trace inner product if ${\bf A}\mathbin{:}{\bf B}=0$.
@@
$\left\{\left({\bf A}\otimes{\bf B}\right)\mathbin{:}{\bf C}\right\}_{\mu}^{\nu}$ $=\left\{{\bf A}\otimes{\bf B}\right\}_{\mu\kappa}^{\nu\lambda}C_{\lambda}^{\kappa}$ Eqn. (B13) $=A_{\mu}^{\nu}B_{\kappa}^{\lambda}C_{\lambda}^{\kappa}$ Eqn. (B10) $=A_{\mu}^{\nu}({\bf B}\mathbin{:}{\bf C})$ Eqn. (B15) $=\left\{{\bf A}\left({\bf B}\mathbin{:}{\bf C}\right)\right\}_{\mu}^{\nu}$
@@
$\left\{\left({\bf A}\,\tilde{\otimes}\,{\bf B}\right)\mathbin{:}{\bf C}\right\}_{\mu}^{\nu}$ $=\left\{{\bf A}\,\tilde{\otimes}\,{\bf B}\right\}_{\mu\kappa}^{\nu\lambda}C_{\lambda}^{\kappa}$ Eqn. (B13) $=A_{\mu}^{\lambda}B_{\kappa}^{\nu}C_{\lambda}^{\kappa}$ Eqn. (45) $=A_{\mu}^{\lambda}\left\{{\bf C}\cdot{\bf B}\right\}_{\lambda}^{\nu}$ Eqn. (B7) $=\left\{{\bf A}{\bf C}{\bf B}\right\}_{\mu}^{\nu}$ Eqn. (B7)
@@ C.1 Hermitian Basis Matrices
$\mbox{$\sigma$}_{\mu}\mathbin{:}\mbox{$\sigma$}_{\nu}$ $=$ $2\delta_{\mu\nu}$ (C4)
@@ C.2 Basis Transformations
In general, a unitary matrix has a determinant equal to $\exp(i\phi)$, and multiplication by $\exp(-i\phi/2)$ yields a unimodular matrix (with $|{\bf R}|=1$). Basis transformations defined by unitary matrices are used extensively throughout both Section V and the following sections of the appendix.
@@ C.3 Spectral Decomposition
Both Hermitian and unitary matrices are normal, and each takes diagonal form in the natural basis defined by its eigenvectors. Consequently, any Hermitian matrix can be diagonalized such that it is equivalent to a differential gain transformation (see Section V.2.1) in its natural basis. Similarly, any unitary matrix is equivalent to differential phase (see Section V.2.2) in its natural basis. This is applied in Proof V.2, where rotation about the line of sight is shown to be equivalent to differential phase between the circularly-polarized natural modes of a magnetized cold plasma.